OpenStudy (anonymous):

How much interest will be paid in 1 year for a loan of $3600 at 8 1/2% simple interest? $______

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

@puppylover123

TheSmartOne (thesmartone):

8 1/2% is othing but 8.5%

TheSmartOne (thesmartone):

nothing*

TheSmartOne (thesmartone):

\(\Huge Interest=\frac{p \times r \times t}{100}\)

OpenStudy (anonymous):

so i do t*r*t?

OpenStudy (anonymous):

306 i believe

OpenStudy (confluxepic):

Direct answer.

OpenStudy (anonymous):

you mean $306?

OpenStudy (anonymous):

yes

OpenStudy (confluxepic):

http://prntscr.com/5a5kse

TheSmartOne (thesmartone):

So